### v1

@@ Abstract
Text-to-image diffusion models have shown great success in generating high-quality text-guided images. Yet, these models may still fail to semantically align generated images with the provided text prompts, leading to problems like incorrect attribute binding and/or catastrophic object neglect. Given the pervasive object-oriented structure underlying text prompts, we introduce a novel object-conditioned Energy-Based Attention Map Alignment (EBAMA) method to address the aforementioned problems. We show that an object-centric attribute binding loss naturally emerges by approximately maximizing the log-likelihood of a $z$-parameterized energy-based model with the help of the negative sampling technique. We further propose an object-centric intensity regularizer to prevent excessive shifts of objects attention towards their attributes. Extensive qualitative and quantitative experiments, including human evaluation, on several challenging benchmarks demonstrate the superior performance of our method over previous strong counterparts. With better aligned attention maps, our approach shows great promise in further enhancing the text-controlled image editing ability of diffusion models.
@@ 1 Introduction
Recently, large-scale text-to-image diffusion models [25, 18, 21, 26, 1, 10] have showcased remarkable capabilities in producing diverse, imaginative, high-resolution visual content based on free-form text prompts. Despite their revolutionary progress, however, these models may not consistently capture and convey the full semantic meaning of the provided text prompts [4, 24]. Some well-known issues include omission, hallucination, or duplication of details [29], semantic leakage of attributes between entities [24], and miscomprehension of intricate textual descriptions [26].
@@
Many previous works have focused on addressing the semantic misalignment issues, particularly concerning multiple-object generation and attribute binding. Composable Diffusion (CD) [15] composes multiple output noises guided by different objects in a text prompt during the generation process. However, this approach often results in a blend of objects, failing to distinctly separate them. Prompt-to-Prompt (PtP) [7] observes a strong correlation between cross-attention maps and the layout of an image. Building on this, Structured Diffusion (StrD) [6] experiments with averaging attention maps generated by different noun phrases for the same queried image latent representation. Yet, a simple average is inadequate for consistently generating images with multiple objects possessing complex attributes. Attend-and-Excite (AnE) [3] proposes a novel approach of maximizing the attention map scores of object tokens by updating the latent at each sampling step. However, we note that artifacts and incorrect attribute binding are likely when AnE maximizes the attention weights of object tokens without any concerns on attributes. In response, SynGen (SG) [23] proposes an attribute-object pair-centric objective, aiming to minimize the distribution distance within the pair while maximizing it from other tokens, based on the assumption that normlized attention maps follow a multinomial distribution. Our findings (see Figs. 1 and 4-6) indicate that this approach still struggles with object neglect due to its pair-centric nature. We argue that multiple-object generation is more critical than attribute binding, as attributes cannot manifest without the presence of objects. Furthermore, in scenarios with multiple objects and no explicit attributes in a prompt, SG is degraded to standard Stable Diffusion Models (SD) [25]. Diverging from these methods, Energy-Based Cross Attention (EBCA) [19] introduces an Energy-Based Model (EBM) framework [28, 31, 30, 32] for queries and keys within cross-attention mechanisms, proposing updates to text embeddings instead of latent noise representations.
@@
A closer look at both the fluctuations of attention intensities and the attention distributions of attribute-object pairs in these methods shed light on the root cause of the misalignment issues. As illustrated in Fig. 1, alignment in attribute-object attention maps (e.g., ‘purple crown’ in SG) encourages attribute binding. However, attention map alignment alone does not guarantee complete semantic alignment, as the intensity levels of object attention maps are crucial in determining the presence of an object in the final image. For example, in the image generated by SG, the crown is notably absent. Conversely, despite successful generation of both objects with strong intensities, AnE binds the attribute ‘purple’ incorrectly to the suitcase resulting from misaligned attention distributions of attribute-object pairs. Motivated by these key observations, we introduce a novel object-conditioned Energy-Based Attention Map Alignment (EBAMA) method to hopefully address both the incorrect attribute binding and the catastrophic object neglect problems in a unified framework. Notably, we show that approximately maximizing the log-likelihood for a $z$-parameterized EBM effectively leads to optimizing an object-centric binding loss, which emphasizes both the object attention map intensity levels and the attribute-object attention map alignment. We further develope an object-centric intensity regularizer to prevent excessive shifts of objects towards their attributes, providing an extra degree of freedom balancing the trade-off between correct attribute binding and the necessary presence of objects.
@@ 2 Related Work
Most large-scale text-to-image diffusion models [25, 18, 21, 26, 1, 10] utilize classifier-free guidance [8] for improved conditional synthesis results. However, due to its strong linguistic and visual priors injected from the training dataset, these models suffer from diverse semantic misalignment issues related to the objects in the provided text prompts and their attribute(s) [4, 26, 24, 29]. Our approach better aligns images with its provided texts by mitigating the issues of object neglect and incorrect attribute binding without fine-tuning the diffusion models or additional training datasets.
@@ 5.3 Human Evaluation
The human evaluation results are shown in Fig. 7. Notably, the object count in DVMP dataset prompts was evenly distributed among $\{1,2,3\}$. In cases with only one object, $\lambda$ likely has no significant effect, leading to similar performance in image-text alignment regardless of the choices of $\lambda$. This observation partly explains why our method performs similarly to that with $\lambda=0$ Consequently, we recommend selecting different intensity weights based on the number of objects in a prompt to optimize our method’s performance. Furthermore, we observe: 1) our method consistently surpasses SD, AnE, and SG aligned with quantitative results in Tab. 1 and Fig. 3; 2) our method shows a more pronounced advantage over other methods on the natural-language ABC-6K dataset. We argue that our object-centric objective, in harmony with the object-oriented patterns prevalent in naturally occurring prompts, exhibits superior efficacy in handling complex real-world, natural-language-based prompts.
@@ 5.4 Ablation Study
We demonstrate the impact of different choices for the intensity weight $\lambda$, which plays a role in enhancing the intensity level. In Fig. 8, we present some representative examples where the model needs to generate multiple objects with certain modifiers. When $\lambda=0.5$, the generation is balanced. However, when $\lambda=0.0$, all images more or less suffer from object neglect, e.g. the camera in (a), the stuffed bear in (b), and the crown in (c). Conversely, when $\lambda=1.0$, artifacts are likely to appear, e.g. the camera in (a), the brown bear in (b), and the apple in (c); attribute binding becomes less effective, e.g. the purple camera in (a), the stuffed bear in (b) and the purple apple in (c).
@@ 5.5 Augmented Attribute Editing
PtP [7] allows local editing with word swap, adding new phrase, or re-weighting by replacing the attention weights of unchanged tokens, or re-weighting the attention maps of target tokens. This heavily relies on the semantic coupling between tokens and their attention maps. In Fig. 9, we categorize the failure cases in PtP shown in the left panel into four situations: (a) ineffective editing with aligned text-to-image generation; (b) ineffective editing with incorrect attribute binding, e.g. the semantic leakage of ‘pink’; (c) ineffective editing with object neglect, e.g. the ‘apple’; and (d) insignificant editing with aligned text-to-image generation, e.g. the property ‘metal’ for the drum. In contrast, our method effectively enhances the semantic distribution of attention maps, allowing PtP with our approach to apply effective and significant local attribute editing to the original images, as demonstrated in the right panel. More implementation details are provided in supplementary material.
@@
Figure 9: Augmented attribute editing with our method. The left two columns demonstrates the attribute editing results from PtP, while the right two columns demonstrates the results from PtP w. ours. Each row shares the same random seed.
6 Limitations
@@ 7 Conclusion
8 Societal Impact and Ethical Concerns
@@
Supplementary Material
Appendix A Object-Conditioned Energy-Based Model
@@ Appendix B Algorithm
Appendix C Implementation Details
@@
C.1 Hyperparameters
In our approach, we utilize a default fixed guidance scale of 7.5. The update step size is selected as $\alpha=20$. We employ a DDIM sampler with a total of 50 steps. The update of the latent variable $z_{t}$ is confined to the first half of the denoising process, which, in this context, corresponds to the initial 25 steps. Further discussion regarding the step size and the updated timesteps is in Appendix F.
C.2 Parser
Following [23], we utilize the spaCy parser [9], specifically employing the transformer-based en_core_web_trf model. Initially, we identify tokens within the prompt that are tagged as either NOUN or PNOUN , thereby constituting our object set. Subsequently, we extract all modifiers within this set based on a predefined set of syntactic dependencies, which include amod, nmod, compound, npadvmod, and conj. Finally, any NOUN or PNOUN that functions as a modifier for other entity-nouns within the object set is excluded.
C.3 Attention Map Extraction
@@
C.4 Augmented Attribute Editing Setup
@@
Appendix D Computational Efficiency
@@
Appendix E External modifiers incorporation
@@
Appendix F Additional Ablation Experiments
@@
We investigate different settings for the step size $\alpha$, as depicted in Fig. 12. When $\alpha$ is set to 1, the step size is too small, leading to insufficient attribute binding and the inability to generate multiple objects effectively. This is evident from the examples of blue walls in (b), a green pole in (c), and the missing crown in (a). In this setting, the performance of (a) is similar to that of SD as shown in the middle panel of Fig I. Conversely, with $\alpha$ set to 40, the step size becomes excessively large, causing an overemphasis on certain attributes, e.g. blue in (a), blue in (b), black in (c) (note that the building behind is also black).
Appendix G Additional Details on Human Evaluation
@@
Appendix H Additional Qualitative Results
We present additional qualitative results comparing our method with others in Figs. I-I. It is observed that AnE struggles with incorrect attribute binding, whereas SG exhibits issues related to object neglect.
Appendix I Additional Explanatory Results
We present more attention intensity levels and distribution maps in Figs. 16-18.
Figure 14: Additional qualitative comparison. Each column shares the same random seed. AnE: incorrect attribute binding, e.g., red apple, purple clock, green car. SG: object neglect, e.g. crown, clock, car.
SD Refer to caption Refer to caption Refer to caption
AnE Refer to caption Refer to caption Refer to caption
SG Refer to caption Refer to caption Refer to caption
Ours Refer to caption Refer to caption Refer to caption
a gray crown and
a purple balloon and
a green chair and
a purple car
Figure 15: Additional qualitative comparison. Each column shares the same random seed. AnE: object neglect, e.g. red hat, zebra; incorrect attribute binding, e.g. gray shirt, red bat. SG: object neglect, e.g. small stuffed bear, crown, helmet.
SD Refer to caption Refer to caption Refer to caption
AnE Refer to caption Refer to caption Refer to caption
SG Refer to caption Refer to caption Refer to caption
Ours Refer to caption Refer to caption Refer to caption
   A brown bear with red hat and scarf and a small stuffed bear.     a blue zebra and
     A boy in a red shirt with a helmet and yellow bat. a spotted crown
Intensity Level “blue”     “apple” Generated Image
SD Refer to caption Refer to caption Refer to caption
AnE Refer to caption Refer to caption Refer to caption
SG Refer to caption Refer to caption Refer to caption
Ours Refer to caption Refer to caption Refer to caption
Figure 16: Additional explanatory results. The text prompt is “a red suitcase and a blue apple”. Each column shares the same random seed.
Intensity Level “purple”     “car” Generated Image
SD Refer to caption Refer to caption Refer to caption
AnE Refer to caption Refer to caption Refer to caption
SG Refer to caption Refer to caption Refer to caption
Ours Refer to caption Refer to caption Refer to caption
Figure 17: Additional explanatory results. The text prompt is “an orange backpack and a purple car”. Each column shares the same random seed.
Intensity Level “sliced”     “tomato” Generated Image
SD Refer to caption Refer to caption Refer to caption
AnE Refer to caption Refer to caption Refer to caption
SG Refer to caption Refer to caption Refer to caption
Ours Refer to caption Refer to caption Refer to caption
Figure 18: Additional explanatory results. The text prompt is “a purple modern camera and a spotted baby dog and a sliced tomato”. Each column shares the same random seed.
Figure 15: Additional qualitative comparison. Each column shares the same random seed. AnE: object neglect, e.g. red hat, zebra; incorrect attribute binding, e.g. gray shirt, red bat. SG: object neglect, e.g. small stuffed bear, crown, helmet.
Figure 14: Additional qualitative comparison. Each column shares the same random seed. AnE: incorrect attribute binding, e.g., red apple, purple clock, green car. SG: object neglect, e.g. crown, clock, car.

### v2

@@ Abstract
Text-to-image diffusion models have shown great success in generating high-quality text-guided images. Yet, these models may still fail to semantically align generated images with the provided text prompts, leading to problems like incorrect attribute binding and/or catastrophic object neglect. Given the pervasive object-oriented structure underlying text prompts, we introduce a novel object-conditioned Energy-Based Attention Map Alignment (EBAMA) method to address the aforementioned problems. We show that an object-centric attribute binding loss naturally emerges by approximately maximizing the log-likelihood of a $z$-parameterized energy-based model with the help of the negative sampling technique. We further propose an object-centric intensity regularizer to prevent excessive shifts of objects attention towards their attributes. Extensive qualitative and quantitative experiments, including human evaluation, on several challenging benchmarks demonstrate the superior performance of our method over previous strong counterparts. With better aligned attention maps, our approach shows great promise in further enhancing the text-controlled image editing ability of diffusion models. Code is available at https://github.com/YasminZhang/EBAMA.
@@ 1 Introduction
Recently, large-scale text-to-image diffusion models [26, 19, 22, 27, 2, 11] have showcased remarkable capabilities in producing diverse, imaginative, high-resolution visual content based on free-form text prompts. Despite their revolutionary progress, however, these models may not consistently capture and convey the full semantic meaning of the provided text prompts [5, 25]. Some well-known issues include omission, hallucination, or duplication of details [30], semantic leakage of attributes between entities [25], and miscomprehension of intricate textual descriptions [27].00footnotetext: Accepted to European Conference on Computer Vision (ECCV) 2024
@@
Many previous works have focused on addressing the semantic misalignment issues, particularly concerning multiple-object generation and attribute binding. Composable Diffusion (CD) [16] composes multiple output noises guided by different objects in a text prompt during the generation process. However, this approach often results in a blend of objects, failing to distinctly separate them. Prompt-to-Prompt (PtP) [8] observes a strong correlation between cross-attention maps and the layout of an image. Building on this, Structured Diffusion (StrD) [7] experiments with averaging attention maps generated by different noun phrases for the same queried image latent representation. Yet, a simple average is inadequate for consistently generating images with multiple objects possessing complex attributes. Attend-and-Excite (AnE) [4] proposes a novel approach of maximizing the attention map scores of object tokens by updating the latent at each sampling step. However, we note that artifacts and incorrect attribute binding are likely when AnE maximizes the attention weights of object tokens without any concerns on attributes. Similarly, A-star (A) [1] aims to minimize the intersection of different objects’ attention maps. In response, SynGen (SG) [24] proposes an attribute-object pair-centric objective, aiming to minimize the distribution distance within the pair while maximizing it from other tokens, based on the assumption that normlized attention maps follow a multinomial distribution. Our findings (see Figs. 1 and 4-6) indicate that this approach still struggles with object neglect due to its pair-centric nature. We argue that multiple-object generation is more critical than attribute binding, as attributes cannot manifest without the presence of objects. Furthermore, in scenarios with multiple objects and no explicit attributes in a prompt, SG is degraded to standard Stable Diffusion Models (SD) [26]. Diverging from these methods, Energy-Based Cross Attention (EBCA) [20] introduces an Energy-Based Model (EBM) framework [29, 32, 31, 33, 34] for queries and keys within cross-attention mechanisms, proposing updates to text embeddings instead of latent noise representations.
@@
A closer look at both the fluctuations of attention intensities and the attention distributions of attribute-object pairs in these methods shed light on the root cause of the misalignment issues. As illustrated in Fig. 1, alignment in attribute-object attention maps (e.g., ‘purple crown’ in SG) encourages attribute binding. However, attention map alignment alone does not guarantee complete semantic alignment, as the intensity levels of object attention maps are crucial in determining the presence of an object in the final image. For example, in the image generated by SG, the crown is notably absent. Conversely, despite successful generation of both objects with strong intensities, AnE binds the attribute ‘purple’ incorrectly to the suitcase resulting from misaligned attention distributions of attribute-object pairs. Motivated by these key observations, we introduce a novel object-conditioned Energy-Based Attention Map Alignment (EBAMA) method to hopefully address both the incorrect attribute binding and the catastrophic object neglect problems in a unified framework. Notably, we show that approximately maximizing the log-likelihood for a $z$-parameterized EBM effectively leads to optimizing an object-centric binding loss, which emphasizes both the object attention map intensity levels and the attribute-object attention map alignment. We further develop an object-centric intensity regularizer to prevent excessive shifts of objects towards their attributes, providing an extra degree of freedom balancing the trade-off between correct attribute binding and the necessary presence of objects.
@@ 2 Related Work
Most large-scale text-to-image diffusion models [26, 19, 22, 27, 2, 11] utilize classifier-free guidance [9] for improved conditional synthesis results. However, due to its strong linguistic and visual priors injected from the training dataset, these models suffer from diverse semantic misalignment issues related to the objects in the provided text prompts and their attribute(s) [5, 27, 25, 30, 37]. Our approach better aligns images with its provided texts by mitigating the issues of object neglect and incorrect attribute binding without fine-tuning the diffusion models or additional training datasets.
@@ 5.3 Human Evaluation
The human evaluation results are shown in Fig. 7. We observe that: 1) our method consistently surpasses SD, AnE, and SG aligned with quantitative results in Tab. 1 and Fig. 3; 2) our method shows a more pronounced advantage over other methods on the natural-language ABC-6K dataset. We argue that our object-centric objective, in harmony with the object-oriented patterns prevalent in naturally occurring prompts, exhibits superior efficacy in handling complex real-world, natural-language-based prompts.
@@ 5.4 Ablation Study
We demonstrate the impact of different choices for the intensity weight $\lambda$, which plays a role in enhancing the intensity level. In Fig. 8, we present some representative examples where the model needs to generate multiple objects with certain modifiers. When $\lambda=0.5$, the generation is balanced. However, when $\lambda=0.0$, all images more or less suffer from object neglect. Conversely, when $\lambda=1.0$, artifacts are likely to appear and attribute binding becomes less effective.
Table 3: Ablations on object-conditioned perspective and energy function choice.
A-A A-O O-O
ours($\lambda=0$) 0.814 0.851 0.823
- obj cond. 0.814 (-0.0%) 0.832 (-2.2%) 0.813 (-1.2%)
- cos sim. 0.812 (-0.2%) 0.846 (-0.6%) 0.817 (-0.7%)
SG 0.767 (-5.8%) 0.830 (-2.5%) 0.811 (-1.5%)
Object-Conditioned Perspective
Besides being able to handle more flexible prompts where no attribute-object pairs exist, our method is object-centric and thus, in the repulsive term, we only calculate the energy of objects and non-modifiers. The way SG handles it is to treat objects and modifiers equivalently when faced with non-modifiers. In the row ‘- obj cond.’ of Tab. 3, we replace the energy of object and non-modifier $f(A_{s},A_{l})$ with the average energy of $f(A_{s},A_{l})$ and $f(A_{s},A_{m})$, where $s,l,m$ represent object, non-modifier, and modifier, respectively. Note that as no modifiers exist in A-A, the results remain the same. In the other datasets, our object-conditioned perspective plays a vital role in the success of our method as the performance significantly decreases.
Energy function Choice
To calculate KL div., SG assumes attention maps follow a multinomial distribution. Yet, cosine similarity does not pose any assumption on the distribution and achieves superior performance. In the row ‘- cos sim.’ of Tab. 3, we replace cosine similarity with the average KL div..
@@ 5.5 Augmented Attribute Editing
Figure 9: Augmented attribute editing with our method. The left two columns demonstrates the attribute editing results from PtP, while the right two columns demonstrates the results from PtP w. ours.
PtP [8] allows local editing with word swap, adding new phrase, or re-weighting by replacing the attention weights of unchanged tokens, or re-weighting the attention maps of target tokens. This heavily relies on the semantic coupling between tokens and their attention maps. In Fig. 9, we categorize the failure cases in PtP shown in the left panel into four situations: (a) ineffective editing with aligned text-to-image generation; (b) ineffective editing with incorrect attribute binding, e.g. the semantic leakage of ‘pink’; (c) ineffective editing with object neglect, e.g. the ‘apple’; and (d) insignificant editing with aligned text-to-image generation, e.g. the property ‘metal’ for the drum. In contrast, our method effectively enhances the semantic distribution of attention maps, allowing PtP with our approach to apply effective and significant local attribute editing to the original images.
@@ 6 Conclusion
Acknowledgements
The work was partially supported by NSF DMS-2015577 and a gift fund from Amazon. We truly thank the three anonymous reviewers for their valuable comments.
Supplementary Material
Appendix A Limitations
Appendix B Societal Impact and Ethical Concerns
@@
Appendix C Object-Conditioned Energy-Based Model
@@ Appendix D Algorithm
Appendix E Implementation Details
@@
E.1 Hyperparameters
In our approach, we utilize a default fixed guidance scale of 7.5. The update step size is selected as $\alpha=20$. We employ a DDIM sampler with a total of 50 steps. The update of the latent variable $z_{t}$ is confined to the first half of the denoising process, which, in this context, corresponds to the initial 25 steps. Further discussion regarding the step size and the updated timesteps is in Appendix H.
E.2 Parser
Following [24], we utilize the spaCy parser [10], specifically employing the transformer based en_core_web_trf model. Initially, we identify tokens within the prompt that are tagged as either NOUN or PNOUN , thereby constituting our object set. Subsequently, we extract all modifiers within this set based on a predefined set of syntactic dependencies, which include amod, nmod, compound, npadvmod, and conj. Finally, any NOUN or PNOUN that functions as a modifier for other entity-nouns within the object set is excluded.
E.3 Attention Map Extraction
@@
E.4 Augmented Attribute Editing Setup
@@
E.5 A-Star Comparison
As A-STAR has not released their official code of implementation, we display their reported numeric results of T-C Sim. in Tab. 4. The table shows that our method demonstrates significantly superior performance when it comes to more complicated datasets. Note that A-A, A-O, and O-O include 0, 1, and 2 attributes in their prompts, respectively.
Table 4: Comparison with A-Star. * Values copied from the A-Star paper.
Method A-A A-O O-O
AnE* 0.80 (-2.4%) 0.82 (-3.5%) 0.81 (-3.6%)
SG 0.77 (-6.1%) 0.83 (-2.4%) 0.81 (-3.6%)
A-STAR* 0.82 (-0.0%) 0.84 (-1.2%) 0.82 (-2.4%)
Ours 0.82 0.85 0.84
Appendix F Computational Efficiency
@@
Appendix G External modifiers incorporation
@@
Appendix H Additional Ablation Experiments
@@
We investigate different settings for the step size $\alpha$, as depicted in Fig. 12. When $\alpha$ is set to 1, the step size is too small, leading to insufficient attribute binding and the inability to generate multiple objects effectively. This is evident from the examples of blue walls in (b), a green pole in (c), and the missing crown in (a). Conversely, with $\alpha$ set to 40, the step size becomes excessively large, causing an overemphasis on certain attributes, e.g. blue in (a), blue in (b), black in (c) (note that the building behind is also black).
Appendix I Additional Details on Human Evaluation